# Aligning Human and LLM Judgments: Insights from EvalAssist on Task-Specific Evaluations and AI-assisted Assessment Strategy Preferences

### v1

@@ 3. EvalAssist
EvalAssist abstracts the llm-as-a-judge evaluation process into a library of parameterize-able evaluators (the criterion being the parameter), allowing the user to focus on criteria definition. This approach acknowledges that developers often use complex external workflows to adjust configurations (e.g., model temperature) and experiment with different models and prompts to generate responses (Desmond et al., 2024). EvalAssist consists of a web-based user experience, an API, and a Python toolkit. The user interface provides users with a convenient way of iteratively testing and refining LLM-as-a-judge criteria, and supports both direct (rubric-based) and pairwise assessment paradigms (Figure 1), the two most prevalent forms of LLM-as-a-judge evaluation available (Kim et al., 2023c; Zheng et al., 2024a)
@@
Users can choose the evaluation method based on task complexity, receive AI judgment explanations, and view metrics like positional bias. Once users are satisfied with their criteria, they can use the Python toolkit to run bulk evaluations with larger data sets by exporting auto-generated JSON definitions of their criteria into predefined notebooks provided in the toolkit. We also allow users to save their test cases and provide a catalog of predefined criteria. A test case in the Example Catalog includes a criteria definition and the data being evaluated. On the landing page, users can choose between Direct Assessments and Pairwise comparisons.
@@ 3.1. Direct Assessment
In this mode, users evaluate outputs based on a single criterion rubric they define. Users can define task-relevant input data through variables in the task-context (Figure 2), such as, for example, the prompt, the article to summarize, or the source data for content-grounded Q&A. The next section (Figure 3) allows them to define their criteria with a title, criteria description, and an arbitrary number of free-form options (b) the LLM evaluator will have to choose from during assessment. As such, the system supports both binary and multi-level scale assessments. In the Evaluator section, users need to select an LLM judge. Our system currently supports four judges: mixtral-8x7b-instruct-v01, llama-3-8b-instruct, llama-3-70b-instruct, and prometheus-8x7b-v2. In the Test Data Section, users enter the outputs they want to evaluate (we call them the responses; note that it is possible to edit the variable name here too so the data being evaluated can be better referenced in the criteria definition) and optionally the result they would expect for each output. After running the evaluation, the system shows the actual results next to the expected results, including agreement, positional bias if present, certainty scores, and an explanation.
Figure 2. Task Context for the Summarization Task. The Task Context is consistent for both Direct Assessment and Pairwise strategies. Users have the option to break down the context into variables, such as the instruction and article, to simplify reference while developing evaluation criteria.
@@
Figure 4. Results for Direct Assessment. Users can select their expected judgments for the output, which are auto-populated based on the criteria they define (i.e., the scale items created when setting the criteria). The results display the LLM evaluator’s judgments, indicating whether there is agreement between the user and the AI, along with explanations for each result.
@@ 3.2. Pairwise Comparison
In the pairwise comparison mode (Figure 3 (a)), EvalAssist compares multiple outputs (minimum two) pairwise against one-another selecting the one that better fits their criteria. The best output is determined by computing the win rate across all pairwise output comparisons. Similar to Direct Assessment, users can provide task-relevant input data through variables, define a criteria, and select an evaluator LLM. However, options don’t need to be added to pairwise comparisons. After evaluation, we display the results next to the expected results (see Figure 3), including the winner, ranking, and agreement with expected ranking. Users can click on each result to see detailed explanations including positional bias, win-rate, and explanations for the comparisons with the other outputs.
@@
3.3. LLM Evaluation
@@
When users select the ”Evaluate” button, their input is sent to the chosen evaluator. Each evaluator is designed to perform either direct assessment or pairwise evaluation. The main external difference between these two lies in how the input criteria are structured. Internally, evaluators operate as a dialog with a target LLM (large language model) using a set of custom prompts specific to that LLM. First, the LLM is prompted to review the evaluation task, considering the context, criteria, and subject of evaluation. Based on this, the LLM generates an open-ended assessment that explains its decision-making process. This step is inspired by Chain-of-Thought (CoT) prompting (Wei et al., 2022), encouraging the LLM to base its final judgement on its initial reasoning. This generated assessment is then added to the dialog history. Afterward, the LLM is asked to make a final judgement. Instead of having the LLM directly generate a final decision, we provide a set of options as potential completions. The system calculates the log probability of each option’s tokens using a forward pass and selects the option with the highest linear probability sequence. This comparison is done by evaluating the first token of each completion, followed by the second, and so on. For direct assessment, these completions are the user-provided option strings, while in pairwise evaluation, the completions are the two responses being compared (Response 1 and Response 2). Although using log probabilities for determining the LLM’s judgement is less efficient than direct generation, it is significantly more reliable. The final explanation of the judgement summarizes the initial assessment. Positional bias is checked by shuffling the order of the options presented to the LLM and verifying the consistency of its final decision. A visual representation of the algorithm is shown in Figure 6.
4. Methodology
@@
The study aimed to explore several key aspects of how practitioners engage with LLMs as evaluators. Specifically, it sought to understand what practitioners prioritize in evaluation criteria when using LLMs as judges and how these priorities differ (RQ1). Additionally, it examined the strategies users employ to refine their criteria in pursuit of human-AI alignment during LLM-as-a-judge evaluations (RQ2). The study also focused on the impact of task-related factors and judge strategies on how practitioners refine criteria (RQ3). Furthermore, the research investigated how task and judge strategy affect user perceptions of the AI judge, including trust, perception of positional bias, perception of explanations, and cognitive load (RQ4). Lastly, the study aimed to identify which assessment strategy users preferred (RQ5).
@@
We used a within-subjects study design, involving 15 participants recruited internally from our organization. Participants were recruited internally from individuals who had previously used EvalAssist and through announcements posted on the organization’s Slack channels. These messages invited employees to participate in a study, with eligibility focused prior experience with model evaluation. Participants were informed about the study’s purpose and provided consent before taking part. Participants were provided with detailed information about the study’s purpose, procedures, and their rights as participants. We collected demographic information, including participants’ roles at the company, education levels, and prior experience with model evaluation. Participants then completed a practice task to familiarize themselves with the experimental interface and procedures.
@@
The experimental tasks involved creating criteria and evaluating ouput based on their criteria using either the pairwise approach or the direct assessment approach for six task contexts based on 3 tasks (Q&A, Email Generation, Summarization) each seen twice, once under the direct assessment condition and once under the pairwise condition. Each task is described in Table 1.
@@
The Direct Assessment task involved designing a single criterion rubric with options to assess whether output complies with the criteria (as shown in Figure 4), whereas pairwise comparison (see Figure 3 involves defining the criteria and comparing pairs) of generated outputs to see which better matches criteria. While both approaches come with strengths and weaknesses, we set out to examine how they influence the evaluation of criteria and users’ interactions differently.
@@
Each task was followed by a short survey assessing participants’ trust in the AI, satisfaction with the criteria, cognitive load, perception of positional bias, and the explanations provided by the AI. Throughout the study, data were logged on the number of evaluations run, final human agreement with the LLM evaluator, and the time taken to run each evaluation. The following questions were rated on a 5-point Likert scale, with 1 indicating “strongly disagree” and 5 indicating “strongly agree”:
@@
    • I trusted the AI evaluator to to judge the responses. (adapted from (Poursabzi-Sangdeh et al., 2021))
    • How confident were you in the model’s judgements/evaluations? (adapted from (Buçinca et al., 2020))
@@
    • I had to work had to accomplish my level of performance
To mitigate order effects, we used partial counterbalancing. Participants were randomly assigned to different task orders to ensure that the sequence of tasks did not systematically bias the results. Participants were asked to reflect on the different tasks they had completed and respond to questions about their preferences between direct assessment and pairwise assessment. They also described how they interacted with two types of tasks: objective versus subjective. After completing the six tasks, participants were asked to select their preferred assessment strategy and explain the reasons for their preference.
@@ 4.1. Tasks and Evaluation Criteria
Tasks and evaluation criteria were chosen to capture various levels of granularity and criteria specificity (i.e., single dimension of ’preference’ vs. document groundedness) and to reflect a diversity of tasks and respective models. Despite the responses coming from various sources (various models, datasets, etc.), all of the responses were reviewed by the coauthors to ensure variability in the responses. Below we list the task descriptions. Task examples can be seen in Table 1.
  • Article Summarization: While the summaries can be judged on cohesiveness, consistency, fluency, relevance (Bavaresco et al., 2024) we asked participants to define criteria based on the single dimension of “preference”, as seen in work by (Chiang and Lee, 2023a; Li et al., 2023; Liu et al., 2023). For the summarization task, we presented the original reference document and corresponding generated outputs to users to be judged from (Fabbri et al., 2021).
@@
  • Q&A multi-turn: The Q&A task involved a context document, a multi-turn conversation, and a final response to the last question in the conversation. The output was generated using retrieval-augmented generation (RAG) (Lewis et al., 2020). RAG evaluations focused on three aspects: answer relevance (is the answer relevant to the query?), context relevance (is the retrieved context relevant to the query?), and groundedness (is the response supported by the context?) (Yu et al., 2024). The Q&A data was selected from existing HR support documents at ABC Company, with AI-generated responses created by multiple models, including (Chung et al., 2022). One of the question-and-answer sets involved a reference document and a customer inquiry about upgrading their business travel to business class with their corporate credit card. Participants were asked to create a criterion to evaluate the faithfulness of each response to the reference document. A response is considered faithful to a document when it only reflects the information expressed in the document (Madsen et al., 2024).
@@
Article Summarization Dress worn by vivien leigh when she played scarlett o’hara in 1939 film gone with the wind has fetched \$ 137,000 at auction . The dress - a jacket and full skirt ensemble - was worn in several key scenes in the 1939 movie , including when scarlett o’hara encounters rhett butler , played by clark gable , and when she gets attacked in the shanty town .
@@
Inclusive Gift Exchange:
Participate in a gift exchange that embraces the spirit of giving across cultures. Bring a wrapped gift (under \$20) or a small item that represents your cultural heritage.
@@
Table 1. Example Outputs for Different Task Types in the Evaluation Study. This table presents example outputs generated for three distinct tasks: Article Summarization, Q&A in an HR Chatbot, and Email Generation.
Type of Change Original Criteria Changed Criteria
Scale: Absolutely, Somewhat, No Scale: Yes, No
Scale: Yes, No Scale: Excellent, Good, Average, Poor
Maybe: Not sure. Maybe: Not sure.
Minor Edit What is the best summary? Which is the best summary?
Table 2. Types of changes made across every evaluation.

### v2

@@ 4. EvalAssist: System Design
EvalAssist abstracts the llm-as-a-judge evaluation process into a library of parameterize-able evaluators (the criterion being the parameter), allowing the user to focus on criteria definition. This approach acknowledges that developers often use complex external workflows to adjust configurations and experiment with different models and prompts to generate responses (Desmond et al., 2024). EvalAssist consists of a web-based user experience, an API, and a Python toolkit. The user interface provides users with a convenient way of iteratively testing and refining LLM-as-a-judge criteria, and supports both direct (rubric-based) and pairwise assessment paradigms (relation-based) (Figure 2), the two most prevalent forms of LLM-as-a-judge evaluation available (Kim et al., 2023; Zheng et al., 2024). EvalAssist is designed to be model-agnostic, allowing any instruction-tuned model to be used as the evaluator. Our system leverages the ability of models to follow instructions and supports four judges: mixtral-8x7b-instruct-v01, llama-3-8b-instruct, llama-3-70b-instruct, and prometheus-8x7b-v2. These models were selected because they represent state-of-the-art open-source instruction-tuned language models available at the time of the study and demonstrated strong performance according to HELM benchmarks (Liang et al., 2022; for Research on Foundation Models, 2024). Additionally, Prometheus was chosen specifically for its notable performance as an evaluator model in prior studies (Kim et al., 2023).
@@
Users can choose the evaluation method based on task complexity, receive AI judgment explanations, and view metrics like positional bias. Once users are satisfied with their criteria, they can use the Python toolkit to run bulk evaluations with larger data sets by exporting auto-generated JSON definitions of their criteria into predefined notebooks provided in the toolkit. We also allow users to save their test cases and provide a catalog of predefined criteria. A test case in the Example Catalog includes a criteria definition and the data being evaluated. On the landing page, users can choose between direct assessments and pairwise comparisons. To be able to refer to artifacts and documents relevant to criteria generation (i.e., the output is relevant to the ‘dialog’ or faithful to the ‘reference document’), users can define task-relevant input data through variables in the task-context (Figure 3), such as, for example, the prompt, the article to summarize, or the source data for content-grounded Q&A. This feature makes it easier to reference these elements when defining criteria in the evaluation form (Figure 4). Outputs evaluated in EvalAssist may originate from prompts that are not visible to participants. Depending on task design, prompts and associated context can either be shown or withheld. This separation between prompt generation and evaluation supports the study of how criteria generalize across unseen prompts and reinforces the importance of distinguishing prompt-specific tuning from evaluation criteria.
@@ 4.1. Direct Assessment
In this mode, users evaluate outputs based on a single criterion rubric they define. The evaluation criteria forms allow them to define their criteria with a title, criteria description, and an arbitrary number of free-form options the AI Evaluator will have to choose from during assessment (Figure 4(b)). As such, the system supports both binary and multi-level scale assessments. In the Evaluator section, users need to select an AI evaluator. In the Test Data Section, users enter the outputs they want to evaluate (we call them the responses; note that it is possible to edit the variable name here too so the data being evaluated can be better referenced in the criteria definition) and optionally the result they would expect for each output. After running the evaluation, the system shows the actual results next to the expected results, including agreement, positional bias if present, certainty scores, and an explanation.
Figure 3. Task Context for the Summarization Task. The Task Context is consistent for both direct assessment and pairwise comparison strategies. Users have the option to break down the context into variables, such as the instruction and article, to simplify reference while developing evaluation criteria.
@@
Figure 5. Results for direct assessment. Users can select their expected judgments for the output, which are auto-populated based on the criteria they define (i.e., the scale items created when setting the criteria). The results display the AI Evaluator’s judgments, indicating whether there is agreement between the user and the AI, along with explanations for each result.
@@ 4.2. Pairwise Comparison
In the pairwise comparison mode (Figure 4 (a)), EvalAssist compares multiple outputs (minimum two) pairwise against one-another selecting the one that better fits the criteria. The best output is determined by computing the win rate across all pairwise output comparisons. Similar to direct assessment, users can provide task-relevant input data through variables, define a criteria, and select an evaluator LLM. However, options don’t need to be added to pairwise comparisons. After evaluation, we display the results next to the expected results (see Figure 6), including the winner, ranking, and agreement with expected ranking. Explanations in pairwise comparison can be seen in Figure 6(b) and are generated as a result of comparing pairs of responses to be evaluated. In total, $\binom{N}{2}$ comparisons are performed in a pairwise manner, where N is the total number of outputs being evaluated. Each pairwise comparison generates an explanation. The outputs are then ranked based on a win rate metric, similar to (Dubois et al., 2024). The explanations presented in Figure 6(b) correspond to the comparisons of each summary against the highlighted row in Figure 6(a). As a result, users are able to click on each result to see detailed explanations including positional bias, win-rate, and explanations for the comparisons with the other outputs.
@@
4.3. Positional Bias
Positional bias occurs when a model consistently favors one option based solely on its position, rather than its content (Li et al., 2024). In evaluation tasks, such as pairwise comparisons, this bias arises if the model disproportionately selects the option in a particular position (e.g., always favoring the first option) regardless of the actual content. In direct assessments, where a response is evaluated based on specific criteria, the order of these options can be shuffled. To test for positional bias, the evaluation is conducted twice with the options presented in different orders. If the outcomes differ between the two evaluations, positional bias is present, as it indicates that the position of the options influenced the evaluation. Conversely, if the outcomes are the same, it suggests that the model’s assessment was not influenced by the position. In EvalAssist, we include a positional bias indicator for both direct and pairwise assessments. This indicator is displayed for each row in the results and is flagged in red text to highlight inconsistent judgments by the AI evaluators.
4.4. Evaluation
@@
When users select the ”Evaluate” button, their input is sent to the chosen evaluator. Each evaluator is designed to perform either direct assessment or pairwise evaluation. The main external difference between these two lies in how the input criteria is structured. Internally, evaluators operate as a dialog with a target AI evaluator using a set of custom prompts specific to that AI Evaluator. The process consists of an assessment prompt, a summarization prompt, and an answer selection prompt for all AI evaluators. However, prometheus-8x7b-v2 uses only a single prompt, as it has already been trained to provide high-quality explanations (Kim et al., 2024b). Figure 7 demonstrates these prompts for the llama-3-8b-instruct and llama-3-70b-instruct AI evaluators. We include the remaining prompts for each AI evaluator in the Appendix. First, the LLM is prompted to review the evaluation task, considering the context, criteria, and subject of evaluation. The LLM generates an open-ended assessment that explains its decision-making process. This step is inspired by Chain-of-Thought (CoT) prompting (Wei et al., 2022), encouraging the LLM to base its final judgment on its initial reasoning. This generated assessment is then added to the dialog history. Afterward, the LLM is asked to make a final judgment. Instead of having the LLM directly generate a final decision, we provide a set of options as potential completions. The system extracts the log probability of each option’s tokens using a forward pass and selects the option with the highest probability. This comparison is done by evaluating the first token of each completion, followed by the second, and so on. For direct assessment, these completions are the user-provided option strings, while in pairwise evaluation, the completions are the two responses being compared (Response 1 and Response 2). Although using log probabilities for determining the LLM’s judgment is less efficient than direct generation, it is significantly more reliable because it eliminates the possibility of a hallucination which would lead to a failed evaluation. We use log probabilities over pre-generated completions as a proxy for model confidence, allowing for token-level diagnostics and detection of overconfident errors such as confidently hallucinated or off-topic content. This approach requires access to model logits, though EvalAssist can also operate in generation-based workflows when logits are unavailable. Prior work has shown that token-level confidence is a useful signal for identifying hallucinations (Quevedo et al., 2025; Kryscinski et al., 2020), and we draw on these insights to enhance the transparency of AI judgments. The summarization prompt results in a summarization of the initial assessment which serves as the explanation presented to the user. Positional bias is checked by shuffling the order of the options presented to the LLM and verifying the consistency of its final decision. A visual representation of the algorithm is shown in Figure 15 in the Appendix. For both pairwise and direct assessment tasks, EvalAssist includes positional bias analysis as an indicator of evaluation robustness. While positional bias is more commonly associated with pairwise tasks, it can also affect direct assessments, where the ordering of rating options (“Accurate”-“Inaccurate” scale) may influence outcomes. EvalAssist supports inversion of these options to detect such order effects, and we expose this information to users as part of the judgment result interface.
Figure 7. The assessment prompt, summarization prompt and answer selection prompt for the llama-3-8b-instruct and llama-3-70b-instruct AI evaluators in the direct assessment context. The assessment prompt is used first, generating an evaluation of the response based on context and evaluation criteria. This assessment is then passed to a summarization prompt, which condenses the findings, and an answer selection prompt, which validates the response against the evaluation criteria.
5. Experimental Design
@@
To better understand how people use EvalAssist to define criteria and evaluate output, we conducted a study exploring several key questions. Specifically, we sought to understand how participants approach the development and refinement of evaluation criteria when using LLMs as judges, focusing on the strategies, characteristics, and evaluative dimensions that emerge in their criteria creation process (RQ1). The study also focused on the impact of task-related factors and judge strategies on how practitioners refine criteria (RQ2). Furthermore, the research investigated how task and judge strategy affect user perceptions of the AI judge, including trust, perception of positional bias, perception of explanations, and cognitive load (RQ3). Lastly, the study aimed to identify which judgment strategy users preferred (RQ4).
@@
We used a within-subjects study design, involving 15 participants recruited internally from our organization. Participants were recruited internally from individuals who had previously used EvalAssist and through announcements posted on the organization’s Slack channels. These messages invited employees to participate in a study, with eligibility focused prior experience with model evaluation. The study followed company policies on user data. Participation was voluntary, with the option to withdraw anytime. Data was de-identified, stored securely, and analyzed anonymously. Participants consented to the use of de-identified data for current and future research, ensuring confidentiality. Participants were provided with detailed information about the study’s purpose, procedures, and their rights as participants. We collected demographic information, including participants’ roles at the company, education levels, and prior experience with model evaluation. Participants then completed a practice task to familiarize themselves with the experimental interface and procedures.
@@
The experimental tasks involved creating criteria and evaluating output based on their criteria using either the pairwise approach or the direct assessment approach for six task contexts based on 3 tasks (Q&A, Email Generation, Summarization) each seen twice, once under the direct assessment condition and once under the pairwise condition. For each task, participants were instructed to stop the evaluation process when they felt satisfied with the criteria they had created. This approach provided them with autonomy and flexibility to determine when their criteria met their own standards or expectations. Each task is described in Table 1.
@@
Direct assessment involved designing a single criterion rubric with options to assess whether output complies with the criteria (as shown in Figure 5), whereas pairwise comparison (see Figure 4) involves defining the criteria and comparing pairs) of generated outputs to see which better matches criteria. While both approaches come with strengths and weaknesses, we set out to examine how they influence the evaluation of criteria and users’ interactions differently.
@@
Each task was followed by a short survey assessing participants’ trust in the AI, satisfaction with the criteria, cognitive load, perception of positional bias, and the explanations provided by the AI. Throughout the study, data were logged on the number of evaluations run, final human agreement with the AI Evaluator, and the time taken to run each evaluation. The following questions were rated on a 5-point Likert scale, with 1 indicating “strongly disagree” and 5 indicating “strongly agree”:
@@
    • I trusted the AI evaluator to judge the responses. (adapted from (Poursabzi-Sangdeh et al., 2021))
    • How confident were you in the model’s judgments/evaluations? (adapted from (Buçinca et al., 2020))
@@
    • I had to work hard to accomplish my level of performance
To mitigate order effects, we used partial counterbalancing. Participants were randomly assigned to different task orders to ensure that the sequence of tasks did not systematically bias the results. Participants were asked to reflect on the different tasks they had completed and respond to questions about their preferences between direct assessment and pairwise assessment. They also described how they interacted with two types of tasks: objective versus subjective. After completing the six tasks, participants had a reflection phase in which they were asked to select their preferred judgment strategy. During the reflection phase, participants reviewed the completed tasks, focusing on differences in judgment strategies and task outcomes. They were asked to identify their preferred strategy and explain the reasoning behind their choice.
@@ 5.1. Tasks and Evaluation Criteria
Tasks and evaluation criteria were chosen to capture various levels of granularity and criteria specificity (i.e., single dimension of ‘preference’ vs. document groundedness) and to reflect a diversity of tasks and respective models. Our goal was to design criteria that reflect real-world tasks practitioners encounter, which naturally encompass a diversity of task types. We intentionally avoided defining the criteria as “ambiguous vs. specific” or “subjective vs. objective” because this framing did not align with the variety of tasks included. Instead, we focused on understanding the differences in how participants approached these tasks, specifically, alignment, user preferences, and the types of evaluations or behaviors they exhibited. Responses were drawn from a variety of sources (models, datasets), and were reviewed by the coauthors to ensure sufficient variability. Below we list the task descriptions. Task examples can be seen in Table 1.
  • Article Summarization: While the summaries can be judged on cohesiveness, consistency, fluency, relevance (Bavaresco et al., 2024) we asked participants to define criteria based on the single dimension of “preference”, as seen in work by (Chiang and yi Lee, 2023; Li et al., 2023; Liu et al., 2023). For the summarization task, we presented the original reference document and corresponding generated outputs to users to be judged from (Fabbri et al., 2021).
@@
  • Q&A multi-turn: The Q&A task involved a context document, a multi-turn conversation, and a final response to the last question in the conversation. The output was generated using retrieval-augmented generation (RAG) (Lewis et al., 2020). RAG evaluations focused on three aspects: answer relevance (is the answer relevant to the query?), context relevance (is the retrieved context relevant to the query?), and groundedness (is the response supported by the context?) (Yu et al., 2024). The Q&A data was selected from existing HR support documents at our company, with AI-generated responses created by multiple models, including (Chung et al., 2022). One of the question-and-answer sets involved a reference document and a customer inquiry about upgrading their business travel to business class with their corporate credit card. Participants were asked to create a criterion to evaluate the faithfulness of each response to the reference document. A response is considered faithful to a document when it only reflects the information expressed in the document (Madsen et al., 2024).
@@
Article Summarization Dress worn by Vivien Leigh when she played Scarlett O’Hara in the 1939 film *Gone with the Wind* has fetched \$137,000 at auction.
The dress—a jacket and full skirt ensemble—was worn in several key scenes in the movie, including when Scarlett encounters Rhett Butler, played by Clark Gable, and when she is attacked in the shanty town.
@@
– Inclusive Gift Exchange: Participate in a gift exchange that embraces the spirit of giving across cultures. Bring a wrapped gift (under \$20) or a small item that represents your cultural heritage.
@@
Table 1. Example outputs generated for three distinct tasks in the evaluation study: Article Summarization, Q&A in an HR Chatbot, and Email Generation.